OpenStudy (anonymous):

To find the height of a pole, a surveyor moves 80 feet away from the base of the pole and then measures the angle of elevation from the ground to the top of the pole to be 63°. What is the height of the pole? Round answer to the nearest foot.

OpenStudy (anonymous):

a. 161 ft c. 41 ft b. 157 ft d. 45 ft

OpenStudy (anonymous):

draw a picture and use a trig function to make it easier. tan 63 = x/80 so x = 80*tan63 = 157 ft.